A gas sample has an original volume of 600 ml when collected at 720 mm and 35°C. If a change is made in the gas temperature which causes the volume of the gas sample to become 520 ml at 1.00 atm, what is the new temperature?
Answer #1

The Combined Gas Law.
P1 x V1 x T2 = P2 x V2 x T1.
720mmHg x 600mL x T2=760mmHg x 520mL x 35°C (308K).
T2 = (760 x 520 x 308) ÷ (720 x 600)
Final Temp. (T2) = 121,721,600 ÷ 432,000 = 281.8K
281.8K - 273 = 8.8°C